# Chiral Separation Effect from Holographic QCD

### v1

@@ Abstract
We analyze the chiral separation effect (CSE) in QCD by using the gauge/gravity duality. In QCD, this effect arises from a combination of chiral anomalies and the axial $U(1)$ anomaly. Due to the axial gluon anomaly, the value of the CSE conductivity is not determined by the anomalies of QCD but receives radiative corrections, which leads to nontrivial dependence on temperature and density. To analyze this dependence, we use different variants of V-QCD, a complex holographic model, carefully fitted to QCD data. We find our results for the anomalous CSE conductivity at small chemical potential and nonzero temperature to be in good qualitative agreement with recent results from lattice QCD simulations. We furthermore give predictions for the behavior of the conductivity at finite (vectorial and axial) chemical potentials.
@@ 1 Introduction
Anomalies in field theory leave their imprint on the transport properties of the matter they describe. They can give rise exotic transport phenomena, which may lead to observable effects in real-world systems. Examples of such systems include the Dirac and Weyl semimetals, and perhaps the quark-gluon plasma, generated in heavy-ion collisions Kharzeev:2024zzm . In this latter context, the best known exotic phenomenon is the chiral magnetic effect (CME) Kharzeev:2007jp ; Fukushima:2008xe , where the combination of a magnetic field and chirality imbalance (i.e., difference of number densities of left and right handed quarks) leads to a net current in the direction of the magnetic field. Another related phenomenon is the chiral separation effect (CSE) Metlitski:2005pr ; Son:2009tf , where magnetic field at finite density leads to the generation of a chiral current. Interestingly, charge asymmetries potentially related to these effects have indeed been observed in heavy-ion collisions Belmont:2014lta ; STAR:2015wza . These effects may also be studied by using lattice QCD Buividovich:2009wi ; Yamamoto:2011gk ; Yamamoto:2011ks ; Buividovich:2013hza ; Brandt:2023wgf ; Brandt:2024wlw .
@@
The gauge-gravity duality is a useful tool to study anomalous transport in strongly coupled systems. From the dual perspective, chiral symmetry is enhanced to a higher dimensional gauge symmetry, and the corresponding gauge fields describe the dynamics of the chiral currents. The chiral anomalies are then implemented through Chern-Simons terms of the gauge fields. In such setups, universal values for the anomalous conductivities have been demonstrated to hold quite generally Gynther:2010ed ; Amado:2011zx ; Landsteiner:2011cp ; Gursoy:2014boa ; Gursoy:2014ela ; Grozdanov:2016ala . Be that as it may, the anomalies of the dynamical gauge fields arise as some of the high dimensional gauge fields acquire a mass. In the case of QCD, the massive field is the U(1) axial gauge field. As usual, gauge symmetry does not allow one to add a hard mass term, but the mass can be generated by other means. One possibility, which is the mechanism appearing in gravity duals of QCD Schwarz:2001sf ; Sakai:2004cn ; Casero:2007ae ; Arean:2013tja ; Arean:2016hcs is the Stückelberg mechanism where the mass arises due to coupling of the gauge field to an additional scalar field Jimenez-Alba:2014iia . [^1] In QCD, this field is dual to the $F\wedge F$ operator and the corresponding source is the $\theta$-angle. The radiative corrections to anomalous conductivities in such setups have been studied in Jimenez-Alba:2014iia ; Jimenez-Alba:2015awa ; Gallegos:2018ozs ; Rai:2023nxe . In addition, effects of additional mixed gauge-gravitational anomaly have been studied in Landsteiner:2011iq ; Megias:2013joa . For systematics of hydrodynamics and transport in the presence of anomalies, see e.g., Son:2009tf ; Neiman:2010zi ; Kalaydzhyan:2011vx ; Megias:2013joa ; Bu:2018psl ; Bu:2019mow ; Ammon:2020rvg .
@@ 3 Anomalous conductivities
so that we recover the universal result Son:2009tf ; Neiman:2010zi ; Landsteiner:2011cp ; Gursoy:2014boa ; Gursoy:2014ela ; Grozdanov:2016ala
$\sigma^{U}_{VV}=0\ ,\qquad\sigma^{U}_{VA}=\frac{N_{c}N_{f}}{2\pi^{2}}\mu_{V}\ ,\qquad\sigma^{U}_{V\Omega}=0\ ,\qquad(T\to\infty)\ .$ (31)
@@ 4 CSE results
Moving on to Fig. 3, we see in general the same behaviour for the CSE for choices $1$ and $2$ of the $Z(\lambda)$ potential. Even though we cannot deduce a suppression or enhancement with respect to the universal result, we note an apparent weakening of the temperature (and axial chemical potential) dependence for large values of the vector chemical potential. One feature worth commenting on for choice $1$ of the $Z(\lambda)$ potential is the rather strong dependence on the axial chemical potential; the nonlinear dependence on $\mu_{A}$ overtakes the temperature dependence in this case. The large dissimilarity between choices $1$ and $2$ is apparently due to differences in the normalization of $\mu_{A}$ between the potentials [^9] A direct way to estimate the effect of the chemical potential on the background is to evaluate the charges $n_{A}$ defined in Appendix B. We find that for fixed $\mu_{A}$ the charges for choices $1$ and $3$ are roughly of the same size, while the charge for the choice $2$ is significantly smaller. This is in rough agreement with the observed dependence on $\mu_{A}$ in Figs. 2 and 3.. Recall that there is no obvious way to fix these normalizations (see Appendix C).
@@
Notice also that the coefficients $c_{i}$ defined in Appendix A.1 are much larger for the choice 1 of the $Z(\lambda)$ potential – we remind the reader that these coefficients are fixed in order to obtain a good agreement with lattice values of the pseudoscalar glueball spectrum.

### v2

@@ Abstract
We analyze the chiral separation effect (CSE) in QCD by using the gauge/gravity duality. In QCD, this effect arises from a combination of chiral anomalies and the axial $U(1)$ anomaly. Due to the axial gluon anomaly, the value of the CSE conductivity is not determined by the anomalies of QCD but receives radiative corrections, which leads to nontrivial dependence on temperature and density. To analyze this dependence, we use different variants of the V-QCD, a complex holographic model, carefully fitted to QCD data. We find our results for the anomalous CSE conductivity at small chemical potential and nonzero temperature to be in good qualitative agreement with recent results from lattice QCD simulations. We furthermore give predictions for the behavior of the conductivity at finite (vectorial and axial) chemical potentials.
@@ 1 Introduction
Anomalies in field theory leave their imprint on the transport properties of the matter they describe. They can give rise exotic transport phenomena, which may lead to observable effects in real-world systems. Examples of such systems include the Dirac and Weyl semimetals, and perhaps the quark-gluon plasma, generated in heavy-ion collisions Kharzeev:2024zzm . In this latter context, the best known exotic phenomenon is the chiral magnetic effect (CME) Kharzeev:2007jp ; Fukushima:2008xe , where the combination of a magnetic field and chirality imbalance (i.e., difference of number densities of left and right handed quarks) leads to a net current in the direction of the magnetic field. A related phenomenon is the chiral separation effect (CSE) Metlitski:2005pr ; Son:2009tf , where a magnetic field at finite density leads to the generation of a chiral current. The chiral magnetic wave (CMW) Kharzeev:2010gd ; Burnier:2012ae is then a collective hydrodynamic excitation that can arise from the interplay between vector and axial density fluctuations, coupled together by anomalous transport effects (also in the presence of an external magnetic field). Interestingly, charge asymmetries potentially related to these effects have indeed been observed in heavy-ion collisions Belmont:2014lta ; STAR:2015wza ; STAR:2021mii . These effects may also be studied by using lattice QCD Buividovich:2009wi ; Yamamoto:2011gk ; Yamamoto:2011ks ; Buividovich:2013hza ; Brandt:2023wgf ; Brandt:2024wlw .
@@
The gauge-gravity duality is a useful tool to study anomalous transport in strongly coupled systems. From the dual perspective, chiral symmetry is enhanced to a higher dimensional gauge symmetry, and the corresponding gauge fields describe the dynamics of the chiral currents. The chiral anomalies are then implemented through Chern-Simons terms of the gauge fields. In such setups, universal values for the anomalous conductivities have been demonstrated to hold quite generally Gynther:2010ed ; Amado:2011zx ; Landsteiner:2011cp ; Gursoy:2014boa ; Gursoy:2014ela ; Grozdanov:2016ala . Be that as it may, the anomalies of the dynamical gauge fields arise as some of the higher dimensional gauge fields acquire a mass. In the case of QCD, the massive field is the U(1) axial gauge field. As usual, gauge symmetry does not allow one to add a hard mass term, but the mass can be generated by other means. One possibility, which is the mechanism appearing in gravity duals of QCD Schwarz:2001sf ; Sakai:2004cn ; Casero:2007ae ; Arean:2013tja ; Arean:2016hcs is the Stückelberg mechanism where the mass arises due to coupling of the gauge field to an additional scalar field Jimenez-Alba:2014iia . [^1] In QCD, this field is dual to the $F\wedge F$ operator and the corresponding source is the $\theta$-angle. The radiative corrections to anomalous conductivities in such setups have been studied in Jimenez-Alba:2014iia ; Jimenez-Alba:2015awa ; Gallegos:2018ozs ; Rai:2023nxe . In addition, effects of additional mixed gauge-gravitational anomaly have been studied in Landsteiner:2011iq ; Megias:2013joa . For systematics of hydrodynamics and transport in the presence of anomalies, see e.g., Son:2009tf ; Neiman:2010zi ; Kalaydzhyan:2011vx ; Megias:2013joa ; Bu:2018psl ; Bu:2019mow ; Ammon:2020rvg .
@@ 3 Anomalous conductivities
so that we recover the universal result Son:2009tf ; Neiman:2010zi ; Landsteiner:2011cp ; Gursoy:2014boa ; Gursoy:2014ela ; Grozdanov:2016ala  [^9] To reiterate (see footnote 6), some of these references present the conductivities obtained from the covariant current as opposed to the consistent current, as we do here.
$\sigma^{U}_{VV}=0\ ,\qquad\sigma^{U}_{VA}=\frac{N_{c}N_{f}}{2\pi^{2}}\mu_{V}\ ,\qquad\sigma^{U}_{V\Omega}=0\ ,\qquad(T\to\infty)\ ,$ (31)
@@ 4 CSE results
One might also be worried by the fact that even for potentials 3, for which the normalization of the CSE conductivity is well-defined, the results in Fig. 1 differ from the lattice results Brandt:2023wgf by a largish factor at low temperatures, i.e., in the strongly coupled region. However, we think that this disagreement is also a weak-coupling effect: The integral in the expression for the current in (27) is dominated by the near-horizon region and therefore essentially only depends on the values of the fluctuation of axial gauge field $\delta\hat{A}$ in this region. The magnetic field $B_{A}$ is however evaluated at the boundary, thus probing the value of $\delta\hat{A}$ in the ultimate weak-coupling limit.
The flow of $\delta\hat{A}$ between the horizon and boundary region contains the contribution from the region which is described [^10] One can check numerically that the near-boundary expansion is a good approximation for $a\gtrsim 10$. by the near-boundary expansion given in Appendix C and independent of the temperature. This flow gives rise to a constant contribution in the CSE conductivity, i.e., a weak-coupling contribution that is not expected to be reliably captured by the holographic model.
@@
Moving on to Fig. 3, we see in general the same behaviour for the CSE for choices $1$ and $2$ of the $Z(\lambda)$ potential. Even though we cannot deduce a suppression or enhancement with respect to the universal result, we note an apparent weakening of the temperature (and axial chemical potential) dependence for large values of the vector chemical potential. One feature worth commenting on for choice $1$ of the $Z(\lambda)$ potential is the rather strong dependence on the axial chemical potential; the nonlinear dependence on $\mu_{A}$ overtakes the temperature dependence in this case. The large difference between choices $1$ and $2$ is apparently due to differences in the normalization of $\mu_{A}$ between the potentials [^11] A direct way to estimate the effect of the chemical potential on the background is to evaluate the charges $n_{A}$ defined in Appendix B. We find that for fixed $\mu_{A}$ the charges for choices $1$ and $3$ are roughly of the same size, while the charge for the choice $2$ is significantly smaller. This is in rough agreement with the observed dependence on $\mu_{A}$ in Figs. 2 and 3.. Recall that there is no obvious way to fix these normalizations (see Appendix C). Notice also that the coefficients $c_{i}$ defined in Appendix A.1 are much larger for the choice 1 of the $Z(\lambda)$ potential – we remind the reader that these coefficients are fixed in order to obtain a good agreement with lattice values of the pseudoscalar glueball spectrum.
@@ 5 Conclusions
Our $\mu_{V}$-dependent evaluation of the CSE (Figs. 2 and 3) could be plausibly relevant for heavy-ion experiments at the LHC Wang:2023xhn and at RHIC STAR:2021mii . In particular, the recent isobar run at RHIC has been devoted to finding conclusive evidence of the CME [^12] Observation of such a signal has however so far remained elusive.. Our results could also be relevant for the search of the CMW [^13] In more detail, it has been argued that it should be possible to identify a CMW signal, even in the absence of the CME and CSE observations. This stems from the fact that the CMW can occur in the presence of zero average background axial and vector charge density (which are needed to be non-zero to give rise to the CME and CSE respectively). Indeed, the CMW can be generated in chiral matter by density fluctuations as long as there is an external magnetic field., which, while to a lesser extent than the CME, is also the subject of current experimental investigation CMS:2017pah ; STAR:2022zpv .